OpenStudy (anonymous):

find the coordinate of point p which divides the line segment a=(0,4) to b=(6,8) in a ratio of 1:2

OpenStudy (anonymous):

over 6 up 4 maybe over two up oh dang

OpenStudy (anonymous):

over two up \(\frac{4}{3}\) ?

OpenStudy (anonymous):

ok i probably screwed that up let me think for a second

OpenStudy (anonymous):

well actually maybe that is right what method are you supposed to use ?

OpenStudy (anonymous):

i would try \[(2,5\tfrac{1}{3})\] and see if that works

OpenStudy (anonymous):

I'm supposed to use the directed line segment formula i think

OpenStudy (aum):

x = 0 + 1/3 * (6-0) y = 4 + 1/3 * (8-4)

OpenStudy (anonymous):

i wish i knew that that was maybe @aum will know

OpenStudy (anonymous):

looks like what i wrote had no idea that was a formula!!

OpenStudy (anonymous):

tbh idek what it is ... the answer was right though so thanks(:

OpenStudy (anonymous):

yw

OpenStudy (aum):

They want a point 1/3 of ab from a. The x-coordinate will be 1/3 along the "x distance" from a. The y-coordinate will be 1/3 along the "y distance" from a.

OpenStudy (aum):

 IF AC is 1/3 of AB, then, due to similar triangles: AD = 1/3 * AE = 1/3 * (x2-x1) CD = 1/3 * BE = 1/3 * (y2-y1) x coordinate of C = x1 + 1/3 * (x2-x1) y coordinate of C = y1 + 1/3 * (y2-y1)